OpenStudy (anonymous):

Use an inequality to solve the problem. A car rental company has two rental rates. Rate 1 is $63 per day plus $0.18 per mile. Rate 2 is $126 per day plus $0.09 per mile. If you plan to rent for a week how many miles would you need to drive to pay less by taking Rate 2?

OpenStudy (anonymous):

R1(m) = 63 + 0.18m R2(m) = 126 + 0.09m how many mile till they are equal? 63 + 0.18m = 126 + 0.09m 0.18m - 0.09m = 126 - 63 m(0.18 - 0.9) = 63 m = 63/0.9 m = 70

OpenStudy (anonymous):

So I would have to drive 70 miles in order to same $ with rate 2? THANK YOU SOOOOOO MUCH!!!!!

OpenStudy (anonymous):

*save

OpenStudy (anonymous):

yup, and because rate 2 is increasing at a slower rate per mile, it is the cheaper option for driving more than 70 miles

OpenStudy (anonymous):

Wow, I totally faved you. My new BFF. It finally makes sense. Be patient with me, I'm raelly trying to FULLY understand. lol So if driving more than 70 miles rate 2 is best option. but if less than 70 miles rate 1 would be best?

OpenStudy (anonymous):

actually I am kind of confused maybe its just late for this idk, but where did you get m = 63/0.9? I got 700 insteead of 70? HELP!!

OpenStudy (anonymous):

please!!

OpenStudy (anonymous):

yes, less than 70 miles, rate one is best.

OpenStudy (anonymous):

in the 3rd to last line, i evaluated the parentheses, (0.18 - 0.09) = 0.9

OpenStudy (anonymous):

then divide that amount into both sides to get m by itself.

OpenStudy (anonymous):

oh my!appear to have lost a zero somewhere. you are right, it should be 63/0.09 = 700

OpenStudy (anonymous):

ok so use the 700 & take it from there. less than 700 miles will make rate @ cheaper?

OpenStudy (anonymous):

@ *2